# A Model-Consistent Data-Driven Computational Strategy for PDE Joint Inversion Problems

### v2

@@ Abstract
The task of simultaneously reconstructing multiple physical coefficients in partial differential equations (PDEs) from observed data is ubiquitous in applications. In this work, we propose an integrated data-driven and model-based iterative reconstruction framework for such joint inversion problems where additional data on the unknown coefficients are supplemented for better reconstructions. Our method couples the supplementary data with the PDE model to make the data-driven modeling process consistent with the model-based reconstruction procedure. We characterize the impact of learning uncertainty on the joint inversion results for two typical inverse problems. Numerical evidence is provided to demonstrate the feasibility of using data-driven models to improve the joint inversion of multiple coefficients in PDEs.
@@ 1 Introduction
This type of inverse problem appears in many applications where $f$ and $g$ are physical quantities we are interested in imaging from measured data $u_{h}$. For instance, in diffuse optical tomography [3], $f$ and $g$ could be, respectively, the optical absorption and scattering coefficients of the underlying medium to be probed. In photoacoustic imaging [40, 15, 49, 34, 5, 44], $f$ and $g$ could be respectively the ultrasound speed and the optical absorption coefficient of the medium. In geophysical applications such as seismic imaging [14, 22, 28], $f$ and $g$ could be, respectively, the bulk modulus and the density field of the Earth. We refer interested readers to  [8, 12, 38, 48, 27, 47] and references therein for more examples of such multiple-coefficient inverse problems. In Section 3, we will present two concrete examples of such inverse problems.
@@
In many applications, we face the issue that when $f$ and $g$ are treated as independent variables, the data available in $\{u_{h}\}_{h\in H}$ are not sufficient to reconstruct $f$ and $g$ simultaneously. This is the situation in diffuse optical tomography where boundary current data are not enough to uniquely determine both the absorption and the scattering coefficients of the underlying medium [3]. The inversion process could be unstable even when the available data are sufficient to determine the two coefficients uniquely. Therefore, one must introduce additional information to improve joint inversions.
@@
There are roughly two lines of ideas in the literature for joint inversion. The first type of method introduces additional measurement data to enrich the information content of data $\{u_{h}\}_{h\in H}$. The other measurement could come from either the same physics in the original problem [44] or a different but related physical process that is coupled to the original problem [1, 16, 20, 35] (often called model fusion). The second type of method uses a priori information one has on the unknown $f$ and $g$ to improve the reconstruction. This includes, for instance, structural methods where one assumes that $f$ and $g$ share the same structural features that one can impose in the reconstruction process through regularization; see [14, 21] and references therein.
@@
In some biological applications, $f$ and $g$ are physical coefficients related to each other either because they describe physical properties that are connected or because their values change similarly in response to an environmental change. For instance, in diffuse optical tomography, the optical absorption and scattering coefficients of the tissue to be probed change correlatedly when tumorous cells are developed. In such a case, one can potentially utilize the relation between the coefficients to improve their computational reconstruction. In the rest of this work, we consider the situation where we have a large amount of past data on $f$ and $g$ that allow us to learn some simplified relation, denoted by $\mathcal{N}_{\theta}$, parameterized by $\theta\in\Theta$, between features of $f$ and $g$. We then use this learned relation to help the simultaneous reconstruction of $f$ and $g$. Let $\mathcal{F}(f)$ represent the part of the information on $f$ that would like to link with the same information on $g$. Then the process can be written mathematically as
@@
The operator $\mathcal{F}$ is decided by given a priori knowledge of the physics of the problem. Examples of $\mathcal{F}$ for practical applications include, for instance, the gradient operator for situations where only the gradient of the $f$ and $g$ are expected to be related [14], and the truncated Fourier transform for applications where only some (lower) Fourier modes between $f$ and $g$ are expected to be related. In the simplest case where $\mathcal{F}$ is invertible, $\mathcal{F}^{-1}\circ\mathcal{N}_{\theta}\circ\mathcal{F}$ gives the relation between the original coefficients $f$ and $g$.
@@ 2 Computational framework
The inversion process with data (3) and (4) is a two-step procedure. In the first step, we learn from data (3) a relation $\mathcal{N}_{\theta}$ between $\mathcal{F}(f)$ and $\mathcal{F}(g)$ as given in (2), while in the second step, we use this relation and data (4) to reconstruct the coefficients $f$ and $g$. Our idea is philosophically similar to the idea of learning a regularization functional from historical data, such as that in [25], but is fundamentally different in terms of the details of the implementation, including how the learning process is done and how the learned model is utilized in the joint inversion stage.
@@ 2.1 Model-consistent learning from historical data
where $\widetilde{X}:=\mathcal{F}[X]$ denote space of coefficients $g$ in the $\mathcal{F}$ domain. This model-consistent learning process is summarized in Algorithm 1.
@@
where $K$ is relatively small (as we are not attempting to learn precise relations but only reasonably good approximations to help the joint reconstructions in the next step of this data-driven joint inversion procedure), to be specified later in the numerical experiments.
@@ 2.2 Joint reconstruction from new data
We can now perform joint inversion of $f$ and $g$ from given noisy data $\{u_{h}^{\delta}\}$, $\delta$ indicates the level of random noise in the data, with the learned relation $\widetilde{\mathcal{N}}_{\widehat{\theta}}=\mathcal{F}^{-1}\circ\mathcal{N}_{\widehat{\theta}}\circ\mathcal{F}$. We perform the reconstruction with the classical model-based iterative inversion framework, where we pursue the reconstruction by minimizing the mismatch between measured data and the model predictions. Instead of using the learned relation between $f$ and $g$ as a hard constraint in the inversion process, we use it as a guide for the reconstruction. We start with the initial solution where we force the relation $\widetilde{\mathcal{N}}_{\widehat{\theta}}$ exactly in the reconstruction. That is, we reconstruct $(\widehat{f}_{0},\widehat{g}_{0}:=\widetilde{\mathcal{N}}_{\widehat{\theta}}(\widehat{f}_{0}))$ with $\widehat{f}_{0}$ given by
@@ 3.1 An inverse diffusion problem
In the first case, we consider a simplified inverse problem in quantitative photoacoustic imaging [5]. Let $\Omega\subseteq\mathbb{R}^{d}$ ($d\geq 1$) be a bounded domain representing an optical medium whose absorption and diffusion coefficients are respectively $\sigma(\mathbf{x})$ and $\gamma(\mathbf{x})$. The transport of near infra-red photons inside the medium can be described by the following diffusion equation
@@
This inverse problem is a simplified model of quantitative photoacoustic tomography, a hybrid acoustic-optical imaging modality where ultrasound measurements are used to get internal data (15); see for instance [5] and references therein for more recent overviews in this direction. The same diffusion model, but with additional boundary current data, is also very useful in practical applications such as diffuse optical tomography [3]. It is well-known that there is an equivalence between $\sigma$ and $\gamma$ in the diffusion equation that prevents the unique reconstruction of the pair $(\sigma,\gamma)$ from a given datum $H$ [5]. It has also been shown [5, 6] that $\sigma$ and $\gamma$ can be uniquely reconstructed with two well-selected datasets when $\gamma$ is known on the boundary of the domain. If that is not the case, then uniqueness with two datasets can not hold [44]. In the rest of this work, we consider the case where additional data on the coefficients $\sigma$ and $\gamma$ are available to allow us to learn a relation between $\sigma$ and $\gamma$ that would help the online reconstruction of the coefficients.
@@ 4.1.1 Experiment I: Learning with polynomial model
In the first numerical experiment, we learn the relationship $\mathcal{N}$ between $\sigma$ and $\gamma$ using the polynomial model (10). We consider the diffusion coefficient $\gamma$ and the absorption coefficient $\sigma$ as Gaussian functions. To be precise, the coefficients are in the forms:
@@
with $a_{i,j}$ ($1\leq i,j\leq 5$) the coupling coefficients.
@@
Learning and testing performance. We perform learning using the polynomial model (10) with $n=2$ following the computational procedure documented in section A.1. To demonstrate the success in learning, we present in Figure 2 the prediction of three randomly selected absorption fields $\sigma$ from the testing dataset, which are generated by the diffusion fields $\gamma$ based on (18)-(19). From left to right, we show the true diffusion fields, the corresponding exact absorption fields, the absorption fields predicted by the learned polynomial model, and the errors in the prediction. We see that the learned polynomial model is pretty successful regarding the accuracy of the prediction.
Figure 3: Internal data $H$ generated with illuminating source (20) for Numerical Experiment I. From left to right are respectively noise-free data, data with $5\%$ additive Gaussian noise, and data with $5\%$ multiplication Gaussian noise.
@@
To test the stability of the proposed coupled scheme, we add Gaussian noise with zero mean and $5\%$ standard derivation to the data used. Figure 4 displays the reconstructed diffusion field, as well as the reconstructed absorption fields. We observe that the reconstructions are fairly accurate in general. Though the resolution of the reconstructions from the noisy data is not as good as the one from the noise-free data, the bump locations for both the diffusion and the absorption fields are clear and accurate.
@@ 4.1.2 Experiment II: Learning with neural network model
In Figure 6, we show three randomly selected absorption fields $\sigma$ from the testing dataset, which are generated by the diffusion fields $\gamma$ based on (22), the corresponding neural network predictions, and the errors in the prediction. We also present the training-validation loss curves in Figure 7. To save the presentation space, we only show the training-validation loss curves for the Fourier modes $\widehat{\sigma}_{(0,0)}$, $\widehat{\sigma}_{(3,3)}$, $\widehat{\sigma}_{(5,5)}$, $\widehat{\sigma}_{(1,2)}$, $\widehat{\sigma}_{(3,4)}$, and $\widehat{\sigma}_{(5,3)}$ but we observe very similar curves for other Fourier modes. A quick takeaway from Figure 6 and Figure 7 is that the training process is quite successful as the testing errors are pretty reasonable, especially given that our training dataset is very small, with only $0.8\times 10^{4}$ data points.
@@
In Figure 9, we show the surface plots of the exact and the reconstructed $\gamma$ and $\sigma$ fields from the proposed data-driven joint inversion scheme. Specifically, we display the diffusion fields at the top panel, and the absorption fields are plotted at the bottom panel. From the left to the right are the surface plots of the true coefficients (first column); the reconstructed diffusion/absorption fields with noise-free internal data (second column), with the internal data containing $5\%$ additive Gaussian noise (third column), and with the internal data including $5\%$ multiplication Gaussian noise (fourth column). We observe that the accuracy of the reconstruction of the proposed joint inversion scheme is pretty well. Even with the noisy internal data, we can recover the main features of the diffusion coefficient $\gamma$ and the absorption field $\sigma$.
@@ 4.2 The acoustic inverse problem
We now present some simulation results on the acoustic wave propagation model (16)-(17). We use the transmission geometry where the illuminating sources are placed on one side of the medium while the receivers are placed on the opposite side of the medium. This is the setup in ultrasound transmission tomography [29]; see the setup in the right plot of Figure 1. To mimic wave propagation in the free space, which is the model used in applications, we use the standard technique of perfectly matched layers (PML) [9, 32]. In a nutshell, we surround the medium to be reconstructed with a layer of specially designed (absorbing) medium such that waves exit the medium will not be bounded back into the medium again. The details of the implementation and benchmark of the forward solver are documented in [13].
We use the time-domain stagger-grid finite difference scheme that adopts a second-order in both the time and the spatial directions to solve the wave equation (16), then record the wave signal starting at time $t_{0}=0$ until the termination time $T$ for each time step. The spatial grid is taken to be the uniform Cartesian grid $(x_{i},z_{j})=(i\Delta x,-1+j\Delta z),i,j=0,1,\cdots,M$ with $\Delta x=\Delta z=1/M$. The receivers are equally placed on the bottom boundary, coinciding with the grid points; namely, there are $M+1$ receivers for each pair of $(\kappa,\rho)$. We then record the measured data of the form (17) for the reconstructions.
@@
We perform joint inversions here using the relation (23) in two different numerical simulations. In Figure 11, we show the true coefficients, where $\kappa$ is the superposition of a constant background and a few Gaussian functions, and the reconstructions from data with different noise levels. We plot the wave speed $v:=\sqrt{\frac{\kappa}{\rho}}$ (which is only the definition of wave speed if $\rho$ is a constant) and the density $\rho$. The optimization algorithm for the joint reconstruction is set with the termination tolerances $10^{-8}$ for both the first-order optimality condition and the updating step size. In addition, the hyper-parameters in Algorithm  2 is set to be $(\eta_{j},J)=(0,3)$. Since we used data collected from many illuminations sources, $N_{s}=20$, the effective noise level in the data is likely much lower than the numbers provided. Even so, we observe significant errors in the reconstructions.
We repeat the simulation for a different true $(\kappa,\rho)$ profile where the bulk modulus $\kappa$ is piecewise constant. The results are shown in Figure 12. The algorithmic parameters are the same as in the previous example. The quality of the reconstructions is quite reasonable. Let us emphasize again that the true relation between $\kappa$ and $\rho$ in (23) is used in the joint reconstruction. Our main purpose here is to show that, once we know the relation between $\kappa$ and $\rho$, the joint reconstruction can be done with good quality. We are not here to show that we can learn the relation faithfully (a task that depends on the quality of the historical data of $\kappa$ and $\rho$ as well as the learning algorithm proposed).

### v3

@@ Abstract
The task of simultaneously reconstructing multiple physical coefficients in partial differential equations (PDEs) from observed data is ubiquitous in applications. In this work, we propose an integrated data-driven and model-based iterative reconstruction framework for such joint inversion problems where additional data on the unknown coefficients are supplemented for better reconstructions. Our method couples the supplementary data with the PDE model to make the data-driven modeling process consistent with the model-based reconstruction procedure. This coupling strategy allows us to characterize the impact of learning uncertainty on the joint inversion results for two typical inverse problems. Numerical evidence is provided to demonstrate the feasibility of using data-driven models to improve the joint inversion of multiple coefficients in PDEs.
@@ 1 Introduction
This type of inverse problem appears in many applications where $f$ and $g$ are physical quantities we are interested in imaging from measured data $u_{h}$. For instance, in diffuse optical tomography [4], $f$ and $g$ could be, respectively, the optical absorption and scattering coefficients of the underlying medium to be probed. In photoacoustic imaging [53, 19, 65, 45, 7, 58], $f$ and $g$ could be respectively the ultrasound speed and the optical absorption coefficient of the medium. In geophysical applications such as seismic imaging [18, 27, 34], $f$ and $g$ could be, respectively, the bulk modulus and the density field of the Earth. We refer interested readers to  [3, 10, 14, 15, 50, 63, 33, 62, 71] and references therein for more examples of such multiple-coefficient inverse problems. In Section 3, we will present two concrete examples of such inverse problems.
@@
In many applications, we face the issue that when $f$ and $g$ are treated as independent variables, the data available in $\{u_{h}\}_{h\in H}$ are not sufficient to reconstruct $f$ and $g$ simultaneously. This is the situation in diffuse optical tomography where boundary current data are not enough to uniquely determine both the absorption and scattering coefficients of the underlying medium [4]. The inversion process could be unstable even when the available data are sufficient to uniquely determine the two coefficients. Therefore, one must introduce additional information to improve joint inversions.
@@
There are roughly two lines of ideas in the literature for joint inversion. The first type of method introduces additional measurement data to enrich the information content of the data $\{u_{h}\}_{h\in H}$. The other measurement could come from either the same physics in the original problem [58] or a different but related physical process that is coupled to the original problem [1, 20, 25, 46] (often called model fusion). The second type of method uses a priori information one has on the unknown $f$ and $g$ to improve the reconstruction. This includes, for instance, structural methods where one assumes that $f$ and $g$ share the same structural features that one can impose in the reconstruction process through regularization; see [18, 26] and references therein.
@@
In some biological applications, $f$ and $g$ are physical coefficients related to each other either because they describe physical properties that are connected or because their values change similarly in response to an environmental change. For instance, in diffuse optical tomography, the optical absorption and scattering coefficients of the tissue to be probed change in correlation when tumorous cells are developed. In such a case, one can potentially utilize the relation between the coefficients to improve their computational reconstruction. In the rest of this work, we consider the situation where we have a large amount of past data on $f$ and $g$ that allows us to learn some simplified relation, denoted by $\mathcal{N}_{\theta}$, parameterized by $\theta\in\Theta$, between features of $f$ and $g$. We then use this learned relation to help the simultaneous reconstruction of $f$ and $g$. Let $\mathcal{F}(f)$ represent the part of the information on $f$ that would like to link with the same information on $g$. Then the process can be written mathematically as
@@
The operator $\mathcal{F}$ is decided by the given a priori knowledge of the physics of the problem. Examples of $\mathcal{F}$ for practical applications include, for instance, the gradient operator for situations where only the gradient of the $f$ and $g$ are expected to be related [18], and the truncated Fourier transform for applications where only some (lower) Fourier modes between $f$ and $g$ are expected to be related. In the simplest case where $\mathcal{F}$ is invertible, $\mathcal{F}^{-1}\circ\mathcal{N}_{\theta}\circ\mathcal{F}$ gives the relation between the original coefficients $f$ and $g$.
Using machine learning methods to solve joint inversion problems has been extensively studied in recent years; see, for instance,  [6, 17, 20, 24, 35, 32, 37, 38, 39, 42, 48, 52, 55, 61, 64, 68, 70] and references therein. In most references cited, machine learning methods are used as an inversion method, meaning that the data are already fixed and one is only interested in using machine learning methods to improve or replace a classical model-based inversion scheme. In the current work, we use machine learning as a way to couple two different types of data, that is, the offline historical coefficient data pairs and the online measurement used to infer new coefficient pairs. Through an offline-online coupling scheme, we improve the overall joint reconstruction procedure. Moreover, we will show that our coupling scheme allows us to characterize how the error in the offline learning process is propagated into the online reconstruction of the coefficients with new data.
@@ 2 Computational framework
The inversion process with data (3) and (4) is a two-step procedure. In the first step, we learn from data (3) a relation $\mathcal{N}_{\theta}$ between $\mathcal{F}(f)$ and $\mathcal{F}(g)$ as given in (2). The choice of the features to be learned, that is, the operator $\mathcal{F}$, is of critical importance in this step. More discussions will be provided later in this section. In the second step, we use this relation and data (4) to reconstruct the coefficients $f$ and $g$. Our idea is philosophically similar to the idea of learning a regularization functional from historical data, such as that in [30], but is fundamentally different in terms of the details of the implementation, including how the learning process is done and how the learned model is utilized in the joint inversion stage.
@@ 2.1 Model-consistent learning from historical data
where $\widetilde{X}:=\mathcal{F}[X]$ denotes space of coefficients $g$ in the $\mathcal{F}$ domain. This model-consistent learning process is summarized in Algorithm 1.
@@
where $K$ is relatively small (as we are not attempting to learn precise relations but only reasonably good approximations to help the joint reconstructions in the next step of this data-driven joint inversion procedure), to be specified later in the numerical experiments. Obviously, one does not want to take a $K$ that is too large, as that would make the dimension of the learning problem too big to be practical.
@@ 2.2 Joint reconstruction from new data
We can now perform joint inversion of $f$ and $g$ from given noisy data $\{u_{h}^{\delta}\}$, where $\delta$ indicates the level of random noise in the data, with the learned relation $\widetilde{\mathcal{N}}_{\widehat{\theta}}=\mathcal{F}^{-1}\circ\mathcal{N}_{\widehat{\theta}}\circ\mathcal{F}$. We perform the reconstruction with the classical model-based iterative inversion framework, where we pursue the reconstruction by minimizing the mismatch between measured data and the model predictions. Instead of using the learned relation between $f$ and $g$ as a hard constraint in the inversion process, we use it as a guide for the reconstruction. We start with the initial solution where we force the relation $\widetilde{\mathcal{N}}_{\widehat{\theta}}$ exactly in the reconstruction. That is, we reconstruct $(\widehat{f}_{0},\widehat{g}_{0}:=\widetilde{\mathcal{N}}_{\widehat{\theta}}(\widehat{f}_{0}))$ with $\widehat{f}_{0}$ given by
@@ 3.1 An inverse diffusion problem
In the first case, we consider a simplified inverse problem in quantitative photoacoustic imaging [7]. Let $\Omega\subseteq\mathbb{R}^{d}$ ($d\geq 1$) be a bounded domain representing an optical medium whose absorption and diffusion coefficients are respectively $\sigma(\mathbf{x})$ and $\gamma(\mathbf{x})$. The transport of near-infrared photons inside the medium can be described by the following diffusion equation
@@
This inverse problem is a simplified model of quantitative photoacoustic tomography, a hybrid acoustic-optical imaging modality where ultrasound measurements are used to get internal data (15); see, for instance,  [7] and references therein for more recent overviews in this direction. The same diffusion model, but with additional boundary current data, is also very useful in practical applications such as diffuse optical tomography [4]. It is well-known that there is an equivalence between $\sigma$ and $\gamma$ in the diffusion equation that prevents the unique reconstruction of the pair $(\sigma,\gamma)$ from a given datum $H$ [7]. It has also been shown [7, 8] that $\sigma$ and $\gamma$ can be uniquely reconstructed with two well-selected datasets when $\gamma$ is known on the boundary of the domain. If that is not the case, then uniqueness with two datasets can not hold [58]. In the rest of this work, we consider the case where additional data on the coefficients $\sigma$ and $\gamma$ are available to allow us to learn a relation between $\sigma$ and $\gamma$ that would help the online reconstruction of the coefficients.
@@ 4.1.1 Experiment I: Learning with polynomial model
In the first numerical experiment, we learn the relationship $\mathcal{N}$ between $\sigma$ and $\gamma$ using the polynomial model (10). We consider the diffusion coefficient $\gamma$ and the absorption coefficient $\sigma$ as Gaussian functions. To be precise, the coefficients are in the form:
@@
with $a_{i,j}$ ($1\leq i,j\leq 5$) being the coupling coefficients.
@@
Learning and testing performance. We perform learning using the polynomial model (10) with $n=2$ following the computational procedure documented in Section A.1. To demonstrate the success in learning, we present in Figure 2 the prediction of three randomly selected absorption fields $\sigma$ from the testing dataset, which are generated by the diffusion fields $\gamma$ in (18). From left to right, we show the true diffusion fields, the corresponding exact absorption fields, the absorption fields predicted by the learned polynomial model, and the errors in the prediction. We see that the learned polynomial model gives a reasonable accuracy of the prediction.
Figure 3: Internal data $H$ generated with illuminating source (20) for Numerical Experiment I. From left to right are respectively noise-free data ($H$), data with $5\%$ additive Gaussian noise ($H^{\delta}=H+0.05\mbox{mean}(H)\eta$), and data with $5\%$ multiplication Gaussian noise ($H^{\delta}=H(1+0.05\eta$)). Here, $\eta$ is a standard normal random variable.
@@
To test the stability of the proposed coupled scheme, we add Gaussian noise with zero mean and $5\%$ standard deviation to the data used. Figure 4 displays the reconstructed diffusion field, as well as the reconstructed absorption fields. We observe that the reconstructions are fairly accurate in general. Though the resolution of the reconstructions from the noisy data is not as good as the one from the noise-free data, the bump locations for both the diffusion and the absorption fields are clear and accurate.
@@ 4.1.2 Experiment II: Learning with neural network model
In Figure 6, we show three randomly selected absorption fields $\sigma$ from the testing dataset, which are generated by the diffusion fields $\gamma$ based on (22), the corresponding neural network predictions, and the errors in the prediction. We also present the training-validation loss curves in Figure 7. To save the presentation space, we only show the training-validation loss curves for the Fourier modes $\widehat{\sigma}_{(0,0)}$, $\widehat{\sigma}_{(3,3)}$, $\widehat{\sigma}_{(5,5)}$, $\widehat{\sigma}_{(1,2)}$, $\widehat{\sigma}_{(3,4)}$, and $\widehat{\sigma}_{(5,3)}$ but we observe very similar curves for other Fourier modes. A quick takeaway from Figure 6 and Figure 7 is that the training process is quite successful as the testing errors are pretty reasonable, especially given that our training dataset is very small, with only $0.8\times 10^{4}$ data points. Our interpretation is that the close alignment of the training and validation curves suggests that the model generalizes well to the validation set and indicates that there is no significant overfitting. The large fluctuations observed in both curves are likely due to optimization dynamics. In this work, our goal is to learn an approximate relationship between the parameters rather than an exact one. While it is possible to reduce training and validation errors, as well as potentially mitigate the observed oscillations, by using more advanced optimization algorithms and fine-tuning hyperparameters, we did not pursue this here because our focus remains on capturing a reasonable approximation of the underlying relationship.
@@
In Figure 9, we show the surface plots of the exact and the reconstructed $\gamma$ and $\sigma$ fields from the proposed data-driven joint inversion scheme. Specifically, we display the diffusion fields in the top panel, and the absorption fields are plotted in the bottom panel. From the left to the right are the surface plots of the true coefficients (first column); the reconstructed diffusion/absorption fields with noise-free internal data (second column), with the internal data containing $5\%$ additive Gaussian noise (third column), and with the internal data including $5\%$ multiplication Gaussian noise (fourth column). We observe that the accuracy of the reconstruction of the proposed joint inversion scheme is pretty well. Even with the noisy internal data, we can recover the main features of the diffusion coefficient $\gamma$ and the absorption field $\sigma$.
@@ 4.2 The acoustic inverse problem
We now present some simulation results on the acoustic wave propagation model (16)-(17). We use the transmission geometry where the illuminating sources are placed on one side of the medium, while the receivers are placed on the opposite side of the medium. This is the setup in ultrasound transmission tomography [36]; see the setup in the right plot of Figure 1. To mimic wave propagation in the free space, which is the model used in applications, we use the standard technique of perfectly matched layers (PML) [11, 43]. In a nutshell, we surround the medium to be reconstructed with a layer of specially designed (absorbing) medium such that waves exiting the medium will not be bounced back into the medium again. The details of the implementation and benchmark of the forward solver are documented in [16].
We use the time-domain stagger-grid finite difference scheme that adopts a second-order accuracy in both the time and the spatial directions to solve the wave equation (16), then record the wave signal starting at time $t_{0}=0$ until the termination time $T$ for each time step. The spatial grid is taken to be the uniform Cartesian grid $(x_{i},z_{j})=(i\Delta x,-1+j\Delta z),i,j=0,1,\cdots,M$ with $\Delta x=\Delta z=1/M$. The receivers are equally placed on the bottom boundary, coinciding with the grid points; namely, there are $M+1$ receivers for each pair of $(\kappa,\rho)$. We then record the measured data of the form (17) for the reconstructions.
@@
We perform joint inversions here using the relation (23) in two different numerical simulations. In Figure 11, we show the true coefficients, where $\kappa$ is the superposition of a constant background and a few Gaussian functions, and the reconstructions from data with different noise levels. We plot the wave speed $v:=\sqrt{\frac{\kappa}{\rho}}$ (which is only the definition of wave speed if $\rho$ is a constant) and the density $\rho$. The optimization algorithm for the joint reconstruction is set with the termination tolerances $10^{-8}$ for both the first-order optimality condition and the updating step size. In addition, the hyper-parameters in Algorithm  2 is set to be $(\eta_{j},J)=(0,3)$. Since we used data collected from many illumination sources, $N_{s}=20$, the effective noise level in the data is likely much lower than the numbers provided. Even so, we observe significant errors in the reconstructions.
We repeat the simulation for a different true $(\kappa,\rho)$ profile where the bulk modulus $\kappa$ is piecewise constant. The results are shown in Figure 12. The algorithmic parameters are the same as in the previous example. The quality of the reconstructions is quite reasonable. Let us emphasize again that the true relation between $\kappa$ and $\rho$ in (23) is used in the joint reconstruction. Our main purpose here is to show that, once we know the relation between $\kappa$ and $\rho$, the joint reconstruction can be done with good quality. We are not here to show that we can learn the relation faithfully (a task that depends on the quality of the historical data of $\kappa$ and $\rho$, as well as the learning algorithm proposed).